# Unified Modeling and Structural Optimization of Multi-magnet Embedded Soft Continuum Robots for Enhanced Kinematic Performances

### v1

@@ II-B1 Elastic Potential Energy
In general deformation conditions, the elastic strain can be approximated as a linear function of the joint angle $\boldsymbol{\theta}_{i}$, [15, 28, 26]. The quadratic form of the elastic strain formulates the elastic potential energy as
$\mathcal{E}_{e}$ $=\frac{1}{2}\sum_{i=0}^{N-1}(\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})^{\top}\boldsymbol{\Lambda}_{i}\boldsymbol{(}\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})$
$=\frac{1}{2}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})^{\top}\boldsymbol{\Lambda}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})$
where $\boldsymbol{\Lambda}_{i}=\operatorname{diag}(\frac{2G_{i}I_{i}}{\ell_{i}},\frac{E_{i}I_{i}}{\ell_{i}},\frac{E_{i}I_{i}}{\ell_{i}})$, $G_{i}$, $E_{i}$, $I_{i}$ are the shear modulus, Young’s modulus, and cross-sectional inertia of the $i$th joint, respectively. The block-diagonal matrix $\boldsymbol{\Lambda}$ is composed of $\boldsymbol{\Lambda}_{i}$, referred to as the joint stiffness matrix. We also note $G_{i}=\frac{E_{i}}{2(1+\rho_{i})}$ for the isotropic material used in the MeSCR, where $\rho_{i}$ represents the Poisson’s ratio. The material properties of each joint are determined by whether it comprises a hard-magnetic material or a soft matrix, both of which exhibit a positive Poisson’s ratio.
II-B2 Magnetic Potential Energy
@@
As shown in Fig. 2, each embedded micromagnet is treated as an independent joint, with additional joints added on either side to assist the description. The joint indices of all $N_{m}$ micromagnets constitute an index set $\mathbb{K}\subset\mathbb{Z}$. Given that the robot operates under a constant temperature, the formulation aligns with the standard Helmholtz free energy without thermal variations [10]. Additional terms with thermal effects can be found in [29]. In this case, the magnetic potential energy of the MeSCR is given by
@@
It is worth noting that some operational scenarios of MeSCR (e.g., interaction with human tissue) typically require low-speed motion to ensure operational safety. Therefore, it is generally assumed that MeSCR operates in quasi-static states, where the inertia terms can be neglected. The dynamic analysis of MeSCR is thus reduced to finding the equilibrium between internal/external forces and potential energy, which can be obtained by solving the following simplified Lagrangian equation:
$\nabla_{\boldsymbol{\theta}}\mathcal{E}_{e}+\nabla_{\boldsymbol{\theta}}\mathcal{E}_{m}+\boldsymbol{\tau}(\boldsymbol{\theta})=\mathbf{0}$ (5)
where $\boldsymbol{\tau}(\boldsymbol{\theta})$ is the internal/external applied torques that can be considered (e.g., gravity torques). We further provide the analytical expressions for the gradients and Hessians of the potential energies, which can improve the accuracy and efficiency of calculations. The gradient and Hessian of $\mathcal{E}_{e}$ are straightforward, whereas we focus on computing that of $\mathcal{E}_{m}$.
$\nabla_{\boldsymbol{\theta}_{i}}\mathcal{E}_{m}=$ $-\sum_{k\in\mathbb{K},k\geq i}\left(\left[\mathbf{m}_{k}\right]_{\times}^{\top}\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}\right)^{\top}\mathbf{b}(\mathbf{p}_{k})$ (6)
$-\sum_{k\in\mathbb{K},k\geq i}\left(\frac{\partial\mathbf{b}(\mathbf{p}_{k})}{\partial\mathbf{p}_{k}}\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}\right)^{\top}\mathbf{m}_{k}$
$\triangleq$ $-\underbrace{\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}}_{\begin{subarray}{c}\text{magnetic}\\ \text{torques}\end{subarray}}-\sum_{k\in\mathbb{K},k\geq i}\left(\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}\right)^{\top}\underbrace{\left(\nabla_{\mathbf{p}_{k}}\mathbf{b}\right)\mathbf{m}_{k}}_{\begin{subarray}{c}\text{magnetic}\\ \text{forces}\end{subarray}}$
where $\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}$ represents the right Jacobian of $\mathrm{SO}(3)$ manifold associated with $\boldsymbol{\theta}_{i}$ [30]. We reorganize the expression by grouping the multiplicative terms into matrix $\mathbf{M}_{i}(\boldsymbol{\theta})\in\mathbb{R}^{3\times 3N_{m}}$ and introducing vector $\mathbf{b}=\left[\mathbf{b}_{k_{0}}^{\top},\dots,\mathbf{b}_{k_{N_{m}-1}}^{\top}\right]^{\top}\in\mathbb{B}\subset\mathbb{R}^{3N_{m}}$ to obtain a compact form. For precise calculations, $\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}$ can be derived from (3). However, the magnetic force is usually negligible in comparison to the magnetic torques, which implies the following explicit representation of (5) as
@@
$\underbrace{\left(\boldsymbol{\Lambda}+\frac{\partial^{2}\mathcal{E}_{m}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}^{\top}}+\frac{\partial\boldsymbol{\tau}}{\partial\boldsymbol{\theta}}\right)}_{\triangleq\mathbf{S}}\delta\boldsymbol{\theta}=\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}$
where the Hessian of $\mathcal{E}_{m}$ (denoted in $\mathbf{S}_{m}$) is given in the supplementary material and $\mathbf{S}\in\mathbb{R}^{3N\times 3N}$ denotes the Hessian matrix of stiffness. Hence, the joint-space compliance relation reads
$\delta\boldsymbol{\theta}=\mathbf{S}^{-1}\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}.$ (8)
Premultiplying (8) by the space Jacobian (2) yields the Cartesian compliance mapping from magnetic field perturbations to distal pose variations:
$\delta\mathbf{h}=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{S}^{-1}\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}\triangleq\mathbf{J}_{\mathbf{b}}\delta\mathbf{b}$ (9)
where $\mathbf{J}_{\mathbf{b}}\in\mathbb{R}^{6\times 3N_{m}}$ is referred to as the actuation Jacobian. We close this subsection by introducing two useful propositions:

### v2

@@ II-B Equations of Motion
where $\boldsymbol{u}$ denotes the actuation input. For general deformations, the elastic strain at each joint can be approximated as a linear function of the joint angle $\boldsymbol{\theta}_{i}$ [19, 18]. The quadratic form of the elastic strain yields the elastic potential energy
$\mathcal{E}_{e}=\frac{1}{2}\sum_{i=0}^{N-1}(\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})^{\top}\boldsymbol{\Lambda}_{i}\boldsymbol{(}\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})=\frac{1}{2}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})^{\top}\boldsymbol{\Lambda}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})$
where $\boldsymbol{\Lambda}_{i}=\operatorname{diag}\left(\tfrac{2G_{i}I_{i}}{\ell_{i}},\tfrac{E_{i}I_{i}}{\ell_{i}},\tfrac{E_{i}I_{i}}{\ell_{i}}\right)$, with $G_{i}$, $E_{i}$, and $I_{i}$ denoting the shear modulus, Young’s modulus, and cross-sectional inertia of the $i$-th joint, respectively. The block-diagonal matrix $\boldsymbol{\Lambda}=\operatorname{blkdiag}(\boldsymbol{\Lambda}_{0},\cdots,\boldsymbol{\Lambda}_{N-1})$ is referred to as the stiffness matrix. For isotropic material, $G_{i}=\tfrac{E_{i}}{2(1+\rho_{i})}$, with $\rho_{i}$ being the Poisson’s ratio. The actuation potential energy $\mathcal{E}_{u}(\boldsymbol{\theta},\boldsymbol{u})$ is left in a general form, representing the work done by internal or external actuators. Its gradient with respect to $\boldsymbol{\theta}$ produces the generalized actuation torques $\nabla_{\boldsymbol{\theta}}\mathcal{E}_{u}(\boldsymbol{\theta},\boldsymbol{u})$.
By Lagrangian formulation, neglecting inertial terms in the quasi-static regime, the equilibrium configuration is determined by the equilibrium condition:
$\nabla_{\boldsymbol{\theta}}\mathcal{E}_{e}(\boldsymbol{\theta})+\nabla_{\boldsymbol{\theta}}\mathcal{E}_{u}(\boldsymbol{\theta},\boldsymbol{u})=\boldsymbol{\tau}(\boldsymbol{\theta}),$ (6)
where $\boldsymbol{\tau}(\boldsymbol{\theta})$ denotes additional generalized torques such as gravity or contact forces.
II-C Kinematic Performances
We next introduce the definition of kinematic performance of SCRs with coordinate-invariance. We begin by recalling that a Riemannian manifold is a smooth manifold equipped with an inner product on each tangent space that varies smoothly with the base point. The associated volume form provides a canonical measure for integration. Let $\mathbb{T}\subset\mathbb{R}^{n}$ be the manifold and use the Euclidean metric, then the induced volume form reduces to the standard Lebesgue measure $\Omega_{\mathbb{T}}=\mathrm{d}\boldsymbol{\theta}$.
Let $^{s}_{b}\mathbf{H}:\mathbb{T}\to SE(3)$ be the forward kinematics of the SCR and $\mathbf{J}_{\boldsymbol{\theta}}$ its geometric Jacobian. Equipping $SE(3)$ with the Euclidean metric induces the pullback metric $(\mathbf{J}_{\boldsymbol{\theta}}^{\top}\mathbf{J}_{\boldsymbol{\theta}})$, whose nonzero eigenvalues $\{\lambda_{i}(\mathbf{J}_{\boldsymbol{\theta}}^{\top}\mathbf{J}_{\boldsymbol{\theta}})\}$ characterize the local distortion of $^{s}_{b}\mathbf{H}$. Any symmetric function of these eigenvalues defines a spectral functional $z(\lambda_{1},\ldots,\lambda_{m})$, which serves as a general mathematical form of a kinematic performance measure. A local performance density can then be defined in general as any symmetric spectral functional $z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}))\triangleq z(\lambda_{1}(\boldsymbol{\theta}),\ldots,\lambda_{m}(\boldsymbol{\theta})),$ with typical choices including
$z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}))=\left\{\begin{aligned} &z_{\mathrm{vol}}(\boldsymbol{\theta})=\sqrt{\det(\mathbf{J}_{\boldsymbol{\theta}}\mathbf{J}_{\boldsymbol{\theta}}^{\top})},\\ &z_{\mathrm{dis}}(\boldsymbol{\theta})=\tfrac{1}{2}\operatorname{tr}(\mathbf{J}_{\boldsymbol{\theta}}^{\top}\mathbf{J}_{\boldsymbol{\theta}}),\\ &z_{\kappa}(\boldsymbol{\theta})=\tfrac{\sigma_{\max}(\mathbf{J}_{\boldsymbol{\theta}})}{\sigma_{\min}(\mathbf{J}_{\boldsymbol{\theta}})},\end{aligned}\right$. (7)
corresponding respectively to volume gain, distortion energy, and condition-number-based dexterity [21, 22, 23]. These local measures are extended to global indices by integration with respect to $\Omega_{\mathbb{T}}$ as $\mathcal{Z}\triangleq\int_{\mathbb{T}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}))\ \Omega_{\mathbb{T}}.$ However, since feasible configurations of the SCR are constrained by the equilibrium condition (6), the effective integration domain is the solution set $\mathbb{T}^{*}=\{\boldsymbol{\theta}\in\mathbb{T}\mid\nabla_{\boldsymbol{\theta}}\mathcal{E}(\boldsymbol{\theta},\boldsymbol{u})=\boldsymbol{\tau}(\boldsymbol{\theta})\}$. If the constraint has constant full rank with respect to $\boldsymbol{\theta}$, $\mathbb{T}^{*}$ is a smooth submanifold of $\mathbb{T}$, and the induced Riemannian volume form can be used. More generally, including neighborhoods of singularities or branch points, the equilibrium set $\mathbb{T}^{*}(\boldsymbol{u})$ is countably rectifiable with intrinsic dimension $\dim(\mathbb{T}^{*}(\boldsymbol{u}))$, hence it carries a well-defined Hausdorff measure $\mathcal{H}^{\dim(\mathbb{T}^{*})}$ and supports geometric integration. Consequently, the constrained global performance can be written as an iterated integral over the actuation parameter space,
$\mathcal{Z}=\int_{\mathbb{U}}\bigg(\int_{\mathbb{T}^{*}}z\!\left({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta})\right)\,\mathrm{d}\mathcal{H}^{\dim(\mathbb{T}^{*})}(\boldsymbol{\theta})\bigg)\,\Omega_{\mathbb{U}},$
where $\Omega_{\mathbb{U}}$ denotes the volume measure on the admissible actuation domain $\mathbb{U}$.
III Structural Optimization Methodology
III-A Problem Formulation
Let $\mathbf{a}\in\mathbb{A}$ denote the vector of structural design parameters, such as segment lengths, joint stiffness blocks, or orientations of embedded units. The admissible set $\mathbb{A}$ is specified by manufacturing limits and safety considerations. For each candidate design $\mathbf{a}$, the forward kinematics $^{s}_{b}\mathbf{H}(\boldsymbol{\theta};\mathbf{a})$ and the corresponding Jacobian $\mathbf{J}_{\boldsymbol{\theta}}(\boldsymbol{\theta};\mathbf{a})$ are uniquely determined, and hence the local performance density $z$ and its global counterpart $\mathcal{Z}(\mathbf{a})$ can be evaluated using (7).
The structural optimization problem is then formulated as
$\min_{\mathbf{a}}$ $\mathcal{Z}(\mathbf{a})\;\triangleq\;\int_{\mathbb{T}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta};\mathbf{a}))\,\Omega_{\mathbb{T}}$ (8)
[…]
where $\mathcal{A}(\mathbf{a})\leq 0$ collects algebraic feasibility constraints. Consequently, the global index associated with the structural parameter $(\mathbf{a})$ can be equivalently written as an iterated integral over the actuation domain and the corresponding solution manifold:
$\min_{\mathbf{a}}\mathcal{Z}(\mathbf{a})$ $=\int_{\mathbb{U}}\Bigg(\int_{\mathbb{T}^{*}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta};\mathbf{a}))\,\mathrm{d}\mathcal{H}^{\dim(\mathbb{T}^{*})}(\boldsymbol{\theta})\Bigg)\,\Omega_{\mathbb{U}}.$ (9)
s.t. $\quad\mathcal{A}(\mathbf{a})\leq 0.$
In the following, we discuss two complementary strategies to solve the optimization problem: a simplification under smooth manifold assumptions and a gradient-based optimization strategy for general cases.
III-B Simplification under Smooth Manifold Assumption
If the equilibrium constraint admits a smooth parametrization of joint states by the input space, i.e., $\boldsymbol{\theta}^{*}:\mathbb{U}\to\mathbb{T}$ defines a smooth immersion, then the mapping $\boldsymbol{\theta}^{*}(\boldsymbol{u})$ satisfies the implicit derivative relation:
$\frac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}=-\mathbf{S}(\boldsymbol{\theta},\boldsymbol{u};\mathbf{a})^{-1}\,\mathbf{M}(\boldsymbol{\theta},\boldsymbol{u};\mathbf{a}),$ (10)
where
$\mathbf{S}=\frac{\partial^{2}\mathcal{E}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}}-\frac{\partial\boldsymbol{\tau}}{\partial\boldsymbol{\theta}},\quad\mathbf{M}=\frac{\partial^{2}\mathcal{E}}{\partial\boldsymbol{\theta}\partial\boldsymbol{u}}$ (11)
are the generalized stiffness matrix and characterize the coupling with the input, respectively. The immersion property of $\boldsymbol{\theta}^{*}$ depends on the rank of $\mathbf{S}^{-1}\mathbf{M}$, the mapping is locally smooth and non-degenerate if $\mathbf{S}$ is nonsingular and $\mathbf{M}$ has full column rank. In this case, the feasible set $\mathbb{T}^{*}$ is diffeomorphic to the input domain $\mathbb{U}$, and the inner integral in (8) reduces to an evaluation along the immersion $\boldsymbol{\theta}^{*}(\boldsymbol{u})$. By applying the area formula [24], one obtains
$\int_{\mathbb{T}^{*}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta};\mathbf{a}))\,\mathrm{d}\mathcal{H}^{\dim(\mathbb{T}^{*})}(\boldsymbol{\theta})=z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}^{*};\mathbf{a}))\,\mathcal{J}_{\boldsymbol{\theta}}(\boldsymbol{u};\mathbf{a}),$
where $\mathcal{J}_{\boldsymbol{\theta}}(\boldsymbol{u};\mathbf{a})$ is the Jacobian determinant of the immersion $\boldsymbol{\theta}^{*}(\boldsymbol{u})$, given by
$\mathcal{J}_{\boldsymbol{\theta}}(\boldsymbol{u};\mathbf{a})=\left[\det\left(\frac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}^{\top}\frac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}\right)\right]^{1/2}\triangleq\left[\mbox{det}\left(\mathbf{G}_{\boldsymbol{\theta}}\right)\right]^{1/2}$
where $\mathbf{G}_{\boldsymbol{\theta}}$ denotes the Gram matrix of $\frac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}$. Substituting back, the global optimization problem (8) takes the following reformulation:
$\min_{\mathbf{a}}\mathcal{Z}(\mathbf{a})$ $=\int_{\mathbb{U}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}^{*}(\boldsymbol{u});\mathbf{a}))\,\mathcal{J}_{\boldsymbol{\theta}}(\boldsymbol{u};\mathbf{a})\,\Omega_{\mathbb{U}}$ (12)
$\mathrm{s.t.}$ $\quad\mathcal{A}(\mathbf{a})\leq 0.$
If the structural feasibility condition ($\mathcal{A}(\mathbf{a})\leq 0$) represents bound constraints, it can be treated as part of the admissible design domain $\mathbb{A}$. Within this domain, the optimization reduces to an unconstrained problem, and the optimality is ensured by satisfying the following first-order necessary condition:
$\nabla_{\mathbf{a}}\mathcal{Z}$ $=\int_{\mathbb{U}}\bigg(z\,\nabla_{\mathbf{a}}\mathcal{J}_{\boldsymbol{\theta}}+\mathcal{J}_{\boldsymbol{\theta}}\,\nabla_{\mathbf{a}}z\bigg)\,\Omega_{\mathbb{U}}$
$=\int_{\mathbb{U}}\left[\frac{1}{2}z\mathcal{J}_{\boldsymbol{\theta}}\operatorname{tr}\left(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{\mathbf{a}}\mathbf{G}_{\boldsymbol{\theta}}\right)+\mathcal{J}_{\boldsymbol{\theta}}\nabla_{\mathbf{a}}z\right]\ \Omega_{\mathbb{U}}$
$=\mathbf{0},$
which, under regularity, yields the strong-form pointwise condition
$\bigg(\frac{1}{2}z\,\operatorname{tr}(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{\mathbf{a}}\mathbf{G}_{\boldsymbol{\theta}})+\nabla_{\mathbf{a}}z\bigg)\mathcal{J}_{\boldsymbol{\theta}}=\mathbf{0},\quad\forall\boldsymbol{u}\in\mathbb{U},$ (13)
where $\mathbf{G}_{\boldsymbol{\theta}}$ is the Gram matrix of $\tfrac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}$. Degeneracy occurs when $\mathcal{J}_{\boldsymbol{\theta}}=0$, corresponding to collapsed input directions where actuation can not produce independent generalized torques. If $z$ is a non-zero kinematic performance index, one can further deduce the non-trivial solution of (13) that
$\operatorname{tr}(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{\mathbf{a}}\mathbf{G}_{\boldsymbol{\theta}})=\sum_{i}\frac{\nabla_{\mathbf{a}}\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}{\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}=-2\frac{\nabla_{\mathbf{a}}z}{z}.$ (14)
Geometrically, this condition implies that optimizing the local performance requires structurally modulating the spectrum of the configuration space metric to counteract its distortion.
III-C Gradient-based Optimization for General Cases
The smooth manifold assumption may fail in practice when the generalized stiffness matrix $(\mathbf{S})$ becomes singular or ill-conditioned. In such cases, the implicit relation (10) is nolonger to hold, and the equilibrium mapping $\boldsymbol{\theta}^{*}$ cannot be obtained analytically. To address this difficulty, we reformulate the optimization as a nested energy-minimization problem that unifies both smooth and singular cases within a single computational framework.
III-C1 Inner-level equilibrium computation
For each input $\boldsymbol{u}$ and structural parameter $\mathbf{a}$, the equilibrium configuration is computed as the stationary point of the total potential energy:
$\boldsymbol{\theta}^{*}(\boldsymbol{u};\mathbf{a})=\arg\min_{\boldsymbol{\theta}}\left(\mathcal{E}(\boldsymbol{\theta},\boldsymbol{u};\mathbf{a})-\int\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}\right).$
The minimization is solved iteratively using a trust-region scheme
$(\mathbf{S}+\mu\mathbf{I})\Delta\boldsymbol{\theta}=\boldsymbol{\tau}(\boldsymbol{\theta})-\nabla_{\boldsymbol{\theta}}\mathcal{E},\quad\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}+\Delta\boldsymbol{\theta}.$ (15)
The damping factor $\mu>0$ is adaptively updated based on the ratio of the actual to the predicted energy reductions. During the actuation, the solution for each $\boldsymbol{u}$ is warm-started from the previous equilibrium, enabling smooth continuation across potential branch points.
III-C2 Outer-level structural update
Once the inner equilibrium mapping is obtained, the global performance index is evaluated numerically by quadrature. The structural parameters are then updated following a projected gradient-descent rule:
$\mathbf{a}_{k+1}=\mathcal{P}_{\mathbb{A}_{c}}(\mathbf{a}_{k}-\eta\nabla_{\mathbf{a}}\mathcal{Z}(\mathbf{a}_{k})),\quad\mathbb{A}_{c}=\{\mathbf{a}|\mathcal{A}(\mathbf{a})\leq 0,\mathbf{a}\in\mathbb{A}\}.$
where $\mathcal{P}_{\mathbb{A}_{c}}(\cdot)$ denotes the projection onto the admissible design domain and $\eta>0$ is a step size ensuring monotonic decrease of $\mathcal{Z}$. The gradient $(\nabla_{\mathbf{a}}\mathcal{Z})$ can be approximated by finite differences or obtained via automatic differentiation through the inner minimization loop. Consequently, the complete computational process is summarized in Algorithm 1.
Algorithm 1 Two-level Structural Optimization Framework
1:Initial structural parameters $\mathbf{a}_{0}$, admissible domain $\mathbb{A}_{c}$, actuation set $\mathbb{U}$, initial configurations $\boldsymbol{\theta}_{0}(\boldsymbol{u})$, damping factor $\mu_{0}>0$.
2:Optimized structural parameters $\mathbf{a}^{*}$.
3:for $k=0,1,\ldots$ until convergence do
4:  for each actuation input $\boldsymbol{u}_{j}\in\mathbb{U}$ do
5:    Initialize $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}_{0}(\boldsymbol{u}_{j})$, $\lambda\leftarrow\lambda_{0}$.
6:    repeat
7:     Compute gradient $\nabla_{\boldsymbol{\theta}}\mathcal{E}(\boldsymbol{\theta},\boldsymbol{u}_{j};\mathbf{a}_{k})$
8:     Compute the stiffness matrix $\mathbf{S}$.
9:     Solve and update the trust-region step (15).
10:    until convergence of $\mathcal{E}$
11:    Store equilibrium $\boldsymbol{\theta}^{*}(\boldsymbol{u}_{j};\mathbf{a}_{k})\leftarrow\boldsymbol{\theta}$.
12:  end for
13:  Evaluate the global performance index $\mathcal{Z}(\mathbf{a}_{k})$.
14:  Update structural parameters by projected gradient descent: $\mathbf{a}_{k+1}=\mathcal{P}_{\mathcal{A}_{c}}\!\big(\mathbf{a}_{k}-\eta\,\nabla_{\mathbf{a}}\mathcal{Z}(\mathbf{a}_{k})\big).$
15:end for
16:return $\mathbf{a}^{*}\leftarrow\mathbf{a}_{k}$.
IV Application of the Optimization Framework to Magnetic Actuation
To demonstrate the generality and practical relevance of the proposed optimization framework, it is applied here to MeSCRs, whose actuation arises from the interaction between internal permanent magnets and external magnetic fields.
IV-A Magnetic Actuation
@@
We formulate the magnetic potential energy and its differential properties with respect to the robot configuration. Each embedded magnet is treated as an independent joint, with additional joints added on either side to assist the description. The joint indices of all $N_{m}$ micromagnets constitute an index set $\mathbb{K}\subset\mathbb{Z}$. Given that the robot operates under a constant temperature, the formulation aligns with the standard Helmholtz free energy without thermal variations [25]. In this case, the magnetic potential energy of the MeSCR is given by
@@
The gradient and Hessian of $\mathcal{E}_{e}$ are straightforward, whereas we focus on computing that of $\mathcal{E}_{m}$.
$\nabla_{\boldsymbol{\theta}_{i}}\mathcal{E}_{m}=$ $-\sum_{k\in\mathbb{K},k\geq i}\left(\left[\mathbf{m}_{k}\right]_{\times}^{\top}\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}+\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}(\mathbf{m}_{k}^{\top}\nabla_{\mathbf{p}_{i}})\right)^{\top}\mathbf{b}_{k}$
$\triangleq$ $\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}$
where $\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}$ represents the right Jacobian of $\mathrm{SO}(3)$ manifold associated with $\boldsymbol{\theta}_{i}$ [26]. We reorganize the expression by grouping the multiplicative terms into matrix $\mathbf{M}_{i}(\boldsymbol{\theta})\in\mathbb{R}^{3\times 3N_{m}}$ and introducing vector $\mathbf{b}=\left[\mathbf{b}_{k_{0}}^{\top},\dots,\mathbf{b}_{k_{N_{m}-1}}^{\top}\right]^{\top}\in\mathbb{B}\subset\mathbb{R}^{3N_{m}}$ to obtain a compact form. For precise calculations, $\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}$ can be derived from (4). As a result, the equilibrium condition (6) has the following explicit representation:
@@
$\left(\boldsymbol{\Lambda}+\frac{\partial^{2}\mathcal{E}_{m}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}^{\top}}+\frac{\partial\boldsymbol{\tau}}{\partial\boldsymbol{\theta}}\right)\delta\boldsymbol{\theta}\triangleq\mathbf{S}\delta\boldsymbol{\theta}=\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}$
where $\mathbf{S}\in\mathbb{R}^{3N\times 3N}$ is known as the Hessian matrix of stiffness. Observed that, the magnetic actuation of the MeSCR follows the canonical equilibrium structure of (11), where variations in the magnetic field $\mathbf{b}$ correspond to the control inputs $\boldsymbol{u}$ that drive configuration changes.
IV-B Properties of Magnetic Actuation under Uniform Fields
We further analyze the properties of the actuation mapping in a uniform magnetic field, where the field can be expressed as $\mathbf{b}=\mathbf{U}_{u}\mathbf{b}_{u}$ with $\mathbf{U}_{u}=\mathbf{1}_{N_{m}}\otimes\mathbf{I}_{3}$ and $\mathbf{b}_{u}\in\mathbb{R}^{3}$. We begin by establishing two fundamental properties of the magnetic actuation mapping: